Consider 𝑔(𝑥)=4sin3𝑥

a. Write down the period of 𝑔(𝑥)

b. Write down the number of solutions to the equation 𝑔(𝑥)=3, for 0≤𝑥≤2𝜋

c. Starting with the graph of 𝑦=sin𝑥, state the transformations which can be used to sketch 𝑔(𝑥).


1
Expert's answer
2022-03-18T17:16:31-0400

a)

Let T the main period of the function "y=sin3x", let "f(x)=sin3x", then

"f(x +T)=sin3(x+T)=sin(3x+3T)"

For T to be the period of the function, must be: "sin(3x+3T)=sin3x"

Because the period of "sinx" is "2\pi", then "3T=2\pi\implies T=\frac{2\pi}3"

Answer: "T=\frac{2\pi}3"


b)

"4sin3x=3\implies sin3x=\frac34\implies 3x=arcsin\frac34," and the second solution is "3x=\pi-arcsin\frac34"

Make up a system and add a period "2\pi" :

"\begin{cases}
3x=arcsin\frac34 +2\pi n\\
3x= \pi-arcsin\frac34 + 2\pi n
\end{cases}\implies" "\begin{cases}
x=\frac{arcsin\frac34}3+\frac{2\pi n}3\\
x=\frac{\pi-arcsin\frac34}3+\frac{2\pi n}3
\end{cases}" "n" is integer

for "x\isin[0;2\pi]:"

"n=0\implies x=\frac{arcsin\frac34}3,x=\frac{\pi - arcsin\frac34}3"

"n=1\implies x=\frac{arcsin\frac34}3+\frac{2\pi}3,x=\frac{\pi -arcsin\frac34}3+\frac{2\pi}3"

"n=2\implies x=\frac{arcsin\frac34}3+\frac{4\pi}3,x=\frac{\pi-arcsin\frac34}3+\frac{4\pi}3"

Answer: 6 solutions in "[0;2\pi]"


c)

To make a graph "y=4sin3x" from a graph "y=sinx", it is necessary to first increase the maximum value of "sinx" by 4 times and build "y=4sinx", then reduce the period of the function by 3 times or narrow along OY and build "y=4sin3x"


On picture:

green - "y=sinx"

blue - "y=4sinx"

red - "y=4sin3x"